OpenStudy (anonymous):

. For y = 4x + 2x2 – x3, show how to find the value for y, when x = 2. Show your work

OpenStudy (michele_laino):

please, you have to substitute 2 in place of x, in your polynomial

OpenStudy (anonymous):

so the answer is 23?

OpenStudy (michele_laino):

please you have to calculate this: \[F(2)=4*2+2^{2}-2^{3}=....\]

OpenStudy (michele_laino):

sorry: \[f(2)=4*2+2*2^{2}-2^{3}=..\]

OpenStudy (michele_laino):

please try, keep in mind that 2^2=4 and 2^3=8

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

4x+2x2=x3=40/27

OpenStudy (michele_laino):

please retry

OpenStudy (anonymous):

orry thanks anyway.